OpenStudy (anonymous):

Simplify. Sqrt(3y) over sqrt(y) A- 1/3 B- sqrt(3) C- 3

OpenStudy (astrophysics):

Do you know of the conjugate?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

\[\frac{ \sqrt{3y} }{ y }\times \frac{ \sqrt{y} }{\sqrt{y}}]

OpenStudy (anonymous):

\[\frac{ \sqrt{3y} }{ \sqrt{y} }\times \frac{ \sqrt{y} }{ \sqrt{y} }\]

OpenStudy (anonymous):

do you know how to go from there @marissa101 ?

OpenStudy (anonymous):

Oh ok I got it thx 😃

OpenStudy (anonymous):

welcome :)